# Preference Optimization with Multi-Sample Comparisons

### v1

@@ 5.3 Improving Quality of Creative Fiction Generation
Fiction generation represents a crucial aspect of the broader field of creative writing, and serves as an essential benchmark for evaluating the capabilities of LLMs (Gómez-Rodríguez & Williams, 2023; Mohammadi, 2024). Despite the proficiency of current LLMs in crafting creative narratives, a significant challenge remains in ensuring a diverse representation of genres (Patel et al., 2024; Wang et al., 2024) while maintaining high writing quality. Ideally, when prompted with a consistent fiction topic, LLMs should be capable of generating distinctly different stories across various genres, closely aligning with user intentions, even in the absence of explicit genre specifications in the prompt. This diversity should extend beyond mere lexical variations, incorporating unique genre-specific elements in each narrative output. In this section, we assess whether the proposed mDPO and mIPO can help fine-tune LLMs to achieve higher quality and more diverse fiction generations. Similar to the previous sections, we compare mDPO and mIPO, with their original baselines.
@@
Preference data construction and finetuning. We utilized over 8,000 publically available prompts [^6] https://draftsparks.com/browse/fiction-prompts/ for fiction generation in constructing our preference dataset. Each prompt was submitted five times to both the Llama 2-7B and Llama 3-8B models. For prompts to Llama 3-8B, we explicitly defined the genres including fantasy, sci-fi, mystery, romance, and horror, to enhance genre diversity, while keeping the prompts for Llama 2-7B unchanged. The responses from Llama 3-8B comprised the chosen set in our preference dataset, whereas the responses from Llama 2-7B (Touvron et al., 2023) formed the rejected set. When finetuning with the single-sample DPO and IPO baselines, one response from each chosen and rejected set was selected. Conversely, with mDPO and mIPO, $k$ responses were selected accordingly. We finetuned the Llama 3-8B model using this preference dataset. More details, including specific prompts used during data generation and all the finetuning parameters, are further illustrated in Appendix A.3.1 and A.3.2, respectively.
@@
Evaluation metrics and results. We primarily evaluated the fine-tuned Llama 3-8B using two metrics: writing quality and genre diversity. For assessing writing quality, we employed the evaluation rubrics proposed by Chakrabarty et al. (2024), which measure fiction quality across four dimensions: fluency, flexibility, originality, and elaboration. Each dimension is specifically evaluated using a total of 14 Yes/No questions. To measure the diversity, we measure entropy of the generated genre diversity and lexical diversity, such as distinct-n (Li et al., 2015). In our experiment, we combined all the questions and used GPT-4o (Achiam et al., 2023) as the judge. Each “Yes” was scored as 1 and each “No” as 0, with the maximum possible score for a generated fiction being 14.
@@
Table 3: Quality comparison of creative fiction writing.
@@
We then compared the final scores of Llama 3-8B finetuned with mDPO, mIPO, and their single-sample baselines. Specific evaluation prompts are shown in Appendix A.3.3. We assess genre diversity using both lexical-level metrics, including the number of distinct unigrams (distinct-1) and bigrams (distinct-2) (Li et al., 2015), as well as semantic-level genre distribution identified by GPT-4o. Details regarding the prompt used for genre identification for a given fiction story are provided in Appendix A.3.4.
The results of the fiction writing quality evaluations are presented in Table 3. It is evident that the proposed multi-sample approaches have a clear advantage over the single-sample baselines, despite the more comprehensive and well-rounded challenge of enhancing the creative writing capabilities (Mohammadi, 2024). The diversity evaluation results, based on the number of unique unigrams and bigrams as well as genre distributions, are shown in Table 4 and Fig. 4, respectively. To better illustrate the differences in diversity in fiction generation, we calculate the
@@
KL-divergence between the distributions shown in Fig. 4 and the uniform distribution, with the results presented in the caption. We observe that models finetuned using multi-sample methods exhibit improved diversity at both lexical and semantic levels compared to the baselines.
@@
Figure 7: Evaluation of multi-sample and single-sample comparison under varying label conditions. The left two plots (green) depict performance in a noise-free setting using GPT-4o labels, the middle two plots (red) show results with default (noisy) labeling, and the right three plots (blue) compare the performance gap between noise-free and noisy settings.
@@
Results with varying $k$. The results for varying group sizes $k$ are shown in Fig. 5, where win rates are computed against models trained using DPO and IPO with the same dataset. The results indicate that both mDPO and mIPO significantly outperform the baselines in terms of win rates. Additionally, we perform an ablation study to examine the effect of $k$ using mDPO and mIPO. As illustrated in the leftmost and middle plots of Fig. 6, we observe that as the value of $k$ increases, the win rates improve while the lose rates decrease for most values of $k$, except for $k=5$ for mIPO.

### v2

@@ 5.3 Improving Quality of Creative Fiction Generation
Fiction generation represents a crucial aspect of the broader field of creative writing, and serves as an essential benchmark for evaluating the capabilities of LLMs (Gómez-Rodríguez and Williams, 2023; Mohammadi, 2024). Despite the proficiency of current LLMs in crafting creative narratives, a significant challenge remains in ensuring a diverse representation of genres (Patel et al., 2024; Wang et al., 2024) while maintaining high writing quality. Ideally, when prompted with a consistent fiction topic, LLMs should be capable of generating distinctly different stories across various genres, closely aligning with user intentions, even in the absence of explicit genre specifications in the prompt. This diversity should extend beyond mere lexical variations, incorporating unique genre-specific elements in each narrative output. In this section, we assess whether the proposed mDPO and mIPO can help fine-tune LLMs to achieve higher quality and more diverse fiction generations. Similar to the previous sections, we compare mDPO and mIPO, with their original baselines.
@@
Preference data construction and finetuning. We utilized over 8,000 publically available prompts [^6] https://draftsparks.com/browse/fiction-prompts/ for fiction generation in constructing our preference dataset. Each prompt was submitted five times to both the Llama 2-7B and Llama 3-8B models. For prompts to Llama 3-8B, we explicitly defined the genres including fantasy, sci-fi, mystery, romance, and horror, to enhance genre diversity, while keeping the prompts for Llama 2-7B unchanged. The responses from Llama 3-8B comprised the chosen set in our preference dataset, whereas the responses from Llama 2-7B (Touvron et al., 2023) formed the rejected set. When finetuning with the single-sample DPO and IPO baselines, one response from each chosen and rejected set was selected. Conversely, with mDPO and mIPO, $k$ responses were selected accordingly. We finetuned the Llama 3-8B model using this preference dataset. More details, including finetuning parameters and specific prompts used during data generation, are illustrated in Appendix 7.3.1 and 7.3.2.
Evaluation metrics and results.
@@
We primarily evaluated the fine-tuned Llama 3-8B using two metrics: writing quality and genre diversity. For assessing writing quality, we employed the evaluation rubrics proposed by Chakrabarty et al. (2024), which measure fiction quality across four dimensions: fluency, flexibility, originality, and elaboration. Each dimension is specifically evaluated using a total of 14 Yes/No questions. To measure the diversity, we measure entropy of the generated genre diversity and lexical diversity, such as distinct-n (Li et al., 2015). In our experiment, we combined all the questions and used GPT-4o (Achiam et al., 2023) as the judge. Each “Yes" was scored as 1 and each “No" as 0, with the maximum possible score for a generated fiction being 14. We then compared the final scores of Llama 3-8B finetuned with mDPO, mIPO, and their single-sample baselines.
Specific evaluation prompts are shown in Appendix 7.3.3. We assess genre diversity using both lexical-level metrics, including the number of distinct unigrams (distinct-1) and bigrams (distinct-2) (Li et al., 2015), as well as semantic-level genre distribution identified by GPT-4o. Details regarding the prompt used for genre identification are provided in Appendix 7.3.4.
@@
Table 3: Quality comparison of creative fiction writing. The maximum score a model can achieve is 14.
@@
The results of the fiction writing quality evaluations are presented in Table 3. It is evident that the proposed multi-sample approaches have a clear advantage over the single-sample baselines, despite the more comprehensive and well-rounded challenge of enhancing the creative writing capabilities (Mohammadi, 2024). The diversity evaluation results, based on the number of unique unigrams and bigrams as well as genre distributions, are shown in Table 4 and Fig. 4, respectively.
@@
To better illustrate the differences in diversity in fiction generation, we calculate the KL-divergence between the distributions shown in Fig. 4 and the uniform distribution, with the results presented in the caption. We observe that models finetuned using multi-sample methods exhibit improved diversity at both lexical and semantic levels compared to the baselines.
@@
Figure 7: Evaluation of multi- and single-sample comparison under varying label conditions. The left two plots (green) depict performance in a noise-free setting using GPT-4o labels, the middle two plots (red) show results with default (noisy) labeling, and the right three plots (blue) compare the performance gap between noise-free and noisy settings.
@@
Results with varying $k$. The results for varying group sizes $k$ are shown in Fig. 5, where win rates are computed against models trained using DPO and IPO with the same dataset. The results indicate that both mDPO and mIPO significantly outperform the baselines. Additionally, we perform an ablation study to examine the effect of $k$ using mDPO and mIPO. As illustrated in the leftmost and middle plots of Fig. 6, we observe that as the value of $k$ increases, the win rates improve while the lose rates decrease for most values of $k$, except for $k=5$ for mIPO.